a student deposited her HS graduation money in the amount of $6200 in two different accounts. The annual interest rate on one account was 5.5% and the interest rate on the second was 5.8%. At the end of the first year the interest on the first account was $92.40 more than the interest earned on the second. Determine the amount in each account.
Answer by Boreal(15235)

x=first account
6200-x=second account
.055x-92.40=.058(6200-x)
.055x-92.40=359.6-.058x
0.113x=452
x=$4000@.055=$220
6200-x=$2200@0.058=$127.60, which is $92.40 less than the first.